OpenStudy (anonymous):

Find S25 for the following arithmetic sequence: -7, -2, 3, 8,…

OpenStudy (anonymous):

first find the common difference "d" d = -2-(-7) = -2+7 = 5

OpenStudy (anonymous):

now use this formula to find Sn Sn = a1+(n-1)d here a1 = -7 n = 25 and d = 5 can you do it now???